# Simplifying Debiased Inference via Automatic Differentiation and Probabilistic Programming

### v1

@@ C.1   Theoretical guarantees
C.2   Motivating the plausibility of $\mathcal{R}_{n}=o_{p}(n^{-1/2})$ using von Mises calculus
@@
when the nuisance estimators $\widehat{P}_{1},\widehat{P}_{2},\ldots,\widehat{P}_{k}$ in Algorithm 3 are replaced by perturbations $P_{1,\epsilon},P_{2,\epsilon},\ldots,P_{k,\epsilon}$ of $P$ belonging to smooth submodels $\mathscr{P}(P,\mathcal{M},s_{1}),\mathscr{P}(P,\mathcal{M},s_{2}),\ldots,\mathscr{P}(P,\mathcal{M},s_{k})$, the estimators of the adjoint evaluations $\dot{\theta}_{j,P,\widehat{h}_{\mathrm{pa}(j)}}^{*}(\widehat{f}_{j})$ are replaced by $\delta_{j,\epsilon}$-perturbations of $\dot{\theta}_{j,P_{j,\epsilon},h_{\mathrm{pa}(j),\epsilon}}^{*}(f_{j,\epsilon})$, and the $o_{p}(n^{-1/2})$ term is replaced by $o(\epsilon)$. Concretely, we investigate whether, as $\epsilon\rightarrow 0$, the outputs $h_{k,\epsilon}$ and $f_{0,\epsilon}$ from Algorithm 5 yield a von Mises expansion of the form
@@
Algorithm 5 Modification of Algorithm 3 with nuisance estimates replaced by deterministic perturbations
@@
The $L^{2}(P)$-consistency condition (ii) is a stability requirement that ensures the influence operator output $f_{0,\epsilon}$ of Algorithm 5 converges to its true value $\dot{\psi}_{P}^{*}(1)$ as the perturbations vanish. Before providing a detailed proof, we give a sketch that indicates where (ii) and the compatibility condition are used. Our proof begins by noting that the output $h_{k,\epsilon}$ of the forward pass of Algorithm 5 can be expressed as the evaluation $\phi(\prod_{j=1}^{k}P_{j,\epsilon})$ of a pathwise differentiable parameter $\phi$ defined on the $k$-fold tensor product of the model $\mathcal{M}$, where $\prod_{j=1}^{k}P_{j,\epsilon}$ denotes a product measure. A von Mises expansion of $\phi$ then yields
@@
C.3   Hilbert-valued parameters
@@
or a cross-fitted variant of this estimator. In practice, $M$ can be selected via cross-validation (Luedtke and Chung, 2023). If the nuisances are all compatible and cross-fitting is used, then this estimator is exactly a regularized one-step estimator as presented in Luedtke and Chung (2023), and so the Hilbert-norm convergence guarantees from that work can be immediately applied here. More generally, only minor adaptations of the arguments from that work are required to provide guarantees for this estimator. The outputs of Algorithm 3 can also be used to construct confidence sets for a Hilbert-valued $\psi(P)$ — see Section 5.2 of Luedtke and Chung (2023).
Appendix D Total pathwise differentiability of primitives from Table 1
D.1   Overview
Appendix D.2 presents primitives that depend nontrivially on both of their arguments. Appendices D.3 and D.4 present primitives that depend only on their distribution-valued or Hilbert-valued argument, respectively. For each primitive, we present the form of the primitive and conditions under which it is totally pathwise differentiable. When not specified, $\mathcal{U}$ is taken to be a generic subset of a Hilbert space $\mathcal{T}$ and $\mathcal{W}$ is taken to be a Hilbert space that is a superset of the image of the map $\theta$ defined on ×$\mathcal{M}\times\mathcal{U}$.
@@ D.1   Overview
D.2   Maps depend nontrivially on both of their arguments
D.2.1   Conditional mean
This example is a special case of the $r$-fold conditional mean studied in Appendix D.2.2 with $r=1$. We refer readers there for details.
D.2.2   Multifold conditional mean
@@ D.2.3   Conditional covariance
Fix $s\in\check{\mathcal{M}}_{P}$, $t\in\check{\mathcal{U}}_{u}$, $\{P_{\epsilon}:\epsilon\}\in\mathscr{P}(P,\mathcal{M},s)$, and $\{u_{\epsilon}=(u_{\epsilon,1},u_{\epsilon,2}):\epsilon\}\in\mathscr{P}(u,\mathcal{U},t)$. The total pathwise differentiability of the $\mathcal{U}\rightarrow L^{2}(Q)$ pointwise product $(u_{1},u_{2})\mapsto u_{1}u_{2}$ (Appendix D.4.5) implies that $\{z\mapsto u_{\epsilon,1}(z)u_{\epsilon,2}(z):\epsilon\}\in\mathscr{P}(u_{1}u_{2},L^{2}(Q),t_{1}u_{2}+u_{1}t_{2})$. Moreover, the total pathwise differentiability of coordinate projections (Appendix S2) implies that $\{z\mapsto u_{\epsilon,j}(z):\epsilon\}\in\mathscr{P}(u_{j},L^{2}(Q),t_{j})$, $j\in\{1,2\}$. When combined with the study of the conditional mean operator in Lemma S8, this shows that
@@
where the $o(\epsilon)$ terms converge to zero in $L^{2}(Q_{X})$ faster than $\epsilon$. Finally, applying the pointwise operation $(a,b,c)\mapsto a-bc$ and leveraging results from Appendix D.4.5 shows that $\theta(P_{\epsilon},u_{\epsilon})-\theta(P,u)-\epsilon\,\dot{\theta}_{P,u}(s,t)=o(\epsilon)$, where
@@ 4. Part 4)
$=\left\langle(s,t),(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))\right\rangle_{\dot{\mathcal{M}}_{P}\oplus\dot{\mathcal{U}}_{u}}$
@@ D.2.7   Optimal solution
Compared to Appendix D.2.6, which studied an optimal value map, this appendix requires additional structure on the feasible region of the optimization problem, namely that it be a Euclidean space. It also requires additional differentiability properties on the objective function, such as the second-order differentiability conditions (ii) and (iii), though does not necessarily require others, such as uniform total pathwise differentiability. This mismatch between the conditions for differentiability of the optimal value and solution maps is not overly surprising given that they capture different aspects of the optimization problem.
@@
D.3   Maps that only depend on their distribution-valued argument
D.3.1   General case: pathwise differentiable parameter
@@
In the remainder of Appendix D.3, we establish the pathwise differentiability of maps $\nu:\mathcal{M}\rightarrow\mathcal{W}$. Applying the general results of this Appendix D.3.1 then shows that the parameter $\theta$ with $\theta(P,u)=\nu(P)$ is totally pathwise differentiable at any $(P,u)\in\mathcal{M}\times\mathcal{U}$ with $\dot{\theta}_{P,u}(s,t)=\dot{\nu}_{P}(s)$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P}^{*}(w),0)$, which will verify the claimed form of the adjoint of the differential operator for the pathwise differentiable primitives in Table 1.
D.3.2   Root-density
@@ D.3.3   Conditional density
The map $\nu_{1}$ is defined so that $\nu_{1}(\tilde{P})=(\tilde{p}_{Y|X}/p_{Y|X})^{1/2}$. Lemma S8 in Luedtke and Chung (2023) shows that $\nu_{1}$ is pathwise differentiable at $P$ with $\dot{\nu}_{1,P}(s)(z)=\frac{1}{2}\Pi_{P}(s)(z)$. The adjoint of this map is $\dot{\nu}_{1,P}^{*}(w)(z)=\frac{1}{2}\{w(z)-E_{P}[w(Z)\,|\,X=x]\}$. The map $\zeta_{1}$ is the pointwise operation $\zeta_{1}(u)(z)=u^{2}(z)p_{Y|X}(z)$. By Appendix D.4.5 and the fact that conditional densities of distributions in $\mathcal{M}$ are uniformly bounded, this map is Hadamard differentiable with $\dot{\zeta}_{1,u}(t)=2tup_{Y|X}$ and $\dot{\zeta}_{1,u}^{*}(w)=2wup_{Y|X}$. The map $\zeta_{2}$ is the change of measure studied in Appendix D.4.9 where, in the notation of that appendix, $\lambda_{1}=P$ and $\lambda_{2}=Q$. By results in that appendix, the fact that $\|dQ/dP\|_{L^{\infty}(P)}<\infty$ implies this map is Hadamard differentiable with $\dot{\zeta}_{2,u}(t)=t$ and $\dot{\zeta}_{2,u}^{*}(w)=\frac{dQ}{dP}w$. Theorem 1 from this work (see also Theorem 3.1 in van der Vaart, 1991a) then shows that $\nu=\zeta_{2}\circ\zeta_{1}\circ\nu_{1}$ is Hadamard differentiable with $\dot{\nu}_{P}^{*}(w)=\dot{\nu}_{1,P}^{*}\circ\dot{\zeta}_{1,\nu_{1}(P)}^{*}\circ\dot{\zeta}_{2,\zeta_{1}\circ\nu_{1}(P)}^{*}(w)$, and properties of adjoints of compositions show that $\dot{\nu}_{P}(s)=\dot{\zeta}_{2,\zeta_{1}\circ\nu_{1}(P)}\circ\dot{\zeta}_{1,\nu_{1}(P)}\circ\dot{\nu}_{1,P}(s)$. Plugging in the provided forms of the differential operators and their adjoints gives (S37).
D.3.4   Dose-response function
@@ D.3.5   Counterfactual density
Under causal conditions, this parameter corresponds to the counterfactual density that would be observed if everyone received treatment $A=1$ (Kennedy et al., 2021). Letting $s_{Y\,|\,A,X}$ and $s_{X}$ be as defined in Appendix D.3.4, the results in Luedtke and Chung (2023) show $\nu$ is pathwise differentiable with local parameter
@@
D.4   Maps that only depend on their Hilbert-valued argument
D.4.1   General case: Hadamard differentiable map
@@
In the remainder of Appendix D.4, we establish the Hadamard differentiability of maps $\zeta:\mathcal{U}\rightarrow\mathcal{W}$. Applying the general results of this Appendix D.4.1 then shows that the parameter $\theta$ with $\theta(P,u)=\zeta(u)$ is totally pathwise differentiable at any $(P,u)\in\mathcal{M}\times\mathcal{U}$ with $\dot{\theta}_{P,u}(s,t)=\dot{\zeta}_{u}(t)$ and $\dot{\theta}_{P,u}^{*}(w)=(0,\dot{\zeta}_{u}^{*}(w))$, which will verify the claimed form of the adjoint of the differential operator for the Hadamard differentiable primitives in Table 1.
D.4.2   Inner product
@@
Importantly, the Hilbert space $\mathcal{T}=\mathcal{R}\oplus\mathcal{R}$ is not ambient for this primitive: the evaluation of this primitive inherently depends on the inner product in $\mathcal{R}$. Hence, when used in Algorithm 1, this primitive does not allow $\mathcal{R}$ to depend on the input to the algorithm, $P$. For example, consider the case where $\mathcal{R}=L^{2}(P)$. In this case, a $P$-dependent inner product primitive would take the form $\theta(P,(u_{1},u_{2}))=\langle u_{1},u_{2}\rangle_{L^{2}(P)}$, which depends nontrivially on its distribution-valued argument and therefore should be studied as in Appendix D.2. For the particular case where $\mathcal{R}=L^{2}(P)$, the $P$-dependent inner product primitive is a special case of the multifold mean studied in Appendix D.2.2, and so is totally pathwise differentiable under conditions given there.
D.4.3   Squared norm
@@
Like the inner product primitive in Appendix D.4.2, the Hilbert space $\mathcal{T}$ is not ambient for this primitive: evaluating $\|\cdot\|_{\mathcal{T}}^{2}$ inherently depends on $\mathcal{T}$. Hence, when used in Algorithm 1, this inner product primitive does not allow $\mathcal{T}$ to depend on the input to the algorithm $P$.
D.4.4   Differentiable functions

### v3

@@ B.1   Theoretical guarantees
B.2   Motivating the plausibility of $\mathcal{R}_{n}=o_{p}(n^{-1/2})$ using von Mises calculus
@@
when the nuisance estimators $\widehat{P}_{1},\widehat{P}_{2},\ldots,\widehat{P}_{k}$ in Algorithm 3 are replaced by perturbations $P_{1,\epsilon},P_{2,\epsilon},\ldots,P_{k,\epsilon}$ of $P$ belonging to smooth submodels $\mathscr{P}(P,\mathcal{M},s_{1}),\mathscr{P}(P,\mathcal{M},s_{2}),\ldots,\mathscr{P}(P,\mathcal{M},s_{k})$, the estimators of the adjoint evaluations $\dot{\theta}_{j,P,\widehat{h}_{\mathrm{pa}(j)}}^{*}(\widehat{f}_{j})$ are replaced by $\delta_{j,\epsilon}$-perturbations of $\dot{\theta}_{j,P_{j,\epsilon},h_{\mathrm{pa}(j),\epsilon}}^{*}(f_{j,\epsilon})$, and the $o_{p}(n^{-1/2})$ term is replaced by $o(\epsilon)$. Concretely, we investigate whether, as $\epsilon\rightarrow 0$, the outputs $h_{k,\epsilon}$ and $f_{0,\epsilon}$ from Algorithm S5 yield a von Mises expansion of the form
@@
Algorithm S5 Modification of Algorithm 3 with nuisance estimates replaced by deterministic perturbations
@@
The $L^{2}(P)$-consistency condition (ii) is a stability requirement that ensures the influence operator output $f_{0,\epsilon}$ of Algorithm S5 converges to its true value $\dot{\psi}_{P}^{*}(1)$ as the perturbations vanish. Before providing a detailed proof, we give a sketch that indicates where (ii) and the compatibility condition are used. Our proof begins by noting that the output $h_{k,\epsilon}$ of the forward pass of Algorithm S5 can be expressed as the evaluation $\phi(\prod_{j=1}^{k}P_{j,\epsilon})$ of a pathwise differentiable parameter $\phi$ defined on the $k$-fold tensor product of the model $\mathcal{M}$, where $\prod_{j=1}^{k}P_{j,\epsilon}$ denotes a product measure. A von Mises expansion of $\phi$ then yields
@@
B.3   Hilbert-valued parameters
@@
or a cross-fitted variant of this estimator. In practice, $M$ can be selected via cross-validation (Luedtke and Chung, 2024). If the nuisances are all compatible and cross-fitting is used, then this estimator is exactly a regularized one-step estimator as presented in Luedtke and Chung (2024), and so the Hilbert-norm convergence guarantees from that work can be immediately applied here. More generally, only minor adaptations of the arguments from that work are required to provide guarantees for this estimator. Those guarantees will typically yield a rate of convergence that is slower than $n^{-1/2}$, which occurs because of a bias-variance tradeoff that arises when choosing $M$—see Luedtke and Chung (2024) for details. The outputs of Algorithm 3 can also be used to construct confidence sets for a Hilbert-valued $\psi(P)$ — see Section 5.2 of Luedtke and Chung (2024).
While the above describes a full algorithm for computing $\widehat{\psi}_{M}$, we emphasize that this algorithm is not currently implemented in the proof-of-concept dimple package. We leave its implementation to future work.
Appendix C Total pathwise differentiability of primitives from Table 1
C.1   Overview
Appendix C.2 presents primitives that depend nontrivially on both of their arguments. Appendices C.3 and C.4 present primitives that depend only on their distribution-valued or Hilbert-valued argument, respectively. For each primitive, we present the form of the primitive and conditions under which it is totally pathwise differentiable. When not specified, $\mathcal{U}$ is taken to be a generic subset of a Hilbert space $\mathcal{T}$ and $\mathcal{W}$ is taken to be a Hilbert space that is a superset of the image of the map $\theta$ defined on ×$\mathcal{M}\times\mathcal{U}$.
@@ C.1   Overview
C.2   Maps depend nontrivially on both of their arguments
C.2.1   Conditional mean
This example is a special case of the $r$-fold conditional mean studied in Appendix C.2.2 with $r=1$. We refer readers there for details.
C.2.2   Multifold conditional mean
@@ C.2.3   Conditional covariance
Fix $s\in\check{\mathcal{M}}_{P}$, $t\in\check{\mathcal{U}}_{u}$, $\{P_{\epsilon}:\epsilon\}\in\mathscr{P}(P,\mathcal{M},s)$, and $\{u_{\epsilon}=(u_{\epsilon,1},u_{\epsilon,2}):\epsilon\}\in\mathscr{P}(u,\mathcal{U},t)$. The total pathwise differentiability of the $\mathcal{U}\rightarrow L^{2}(Q)$ pointwise product $(u_{1},u_{2})\mapsto u_{1}u_{2}$ (Appendix C.4.5) implies that $\{z\mapsto u_{\epsilon,1}(z)u_{\epsilon,2}(z):\epsilon\}\in\mathscr{P}(u_{1}u_{2},L^{2}(Q),t_{1}u_{2}+u_{1}t_{2})$. Moreover, the total pathwise differentiability of coordinate projections (Lemma S2) implies that $\{z\mapsto u_{\epsilon,j}(z):\epsilon\}\in\mathscr{P}(u_{j},L^{2}(Q),t_{j})$, $j\in\{1,2\}$. When combined with the study of the conditional mean operator in Lemma S8, this shows that
@@
where the $o(\epsilon)$ terms converge to zero in $L^{2}(Q_{X})$ faster than $\epsilon$. Finally, applying the pointwise operation $(a,b,c)\mapsto a-bc$ and leveraging results from Appendix C.4.5 shows that $\theta(P_{\epsilon},u_{\epsilon})-\theta(P,u)-\epsilon\,\dot{\theta}_{P,u}(s,t)=o(\epsilon)$, where
@@ 4. Part 4)
$=\left\langle(s,t),(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))\right\rangle_{\dot{\mathcal{M}}_{P}\oplus\dot{\mathcal{U}}_{u}}.$
@@ C.2.7   Optimal solution
Compared to Appendix C.2.6, which studied an optimal value map, this appendix requires additional structure on the feasible region of the optimization problem, namely that it be a Euclidean space. It also requires additional differentiability properties on the objective function, such as the second-order differentiability conditions (ii) and (iii), though does not necessarily require others, such as uniform total pathwise differentiability. This mismatch between the conditions for differentiability of the optimal value and solution maps is not overly surprising given that they capture different aspects of the optimization problem.
@@
C.3   Maps that only depend on their distribution-valued argument
C.3.1   General case: pathwise differentiable parameter
@@
In the remainder of Appendix C.3, we establish the pathwise differentiability of maps $\nu:\mathcal{M}\rightarrow\mathcal{W}$. Applying the general results of this Appendix C.3.1 then shows that the parameter $\theta$ with $\theta(P,u)=\nu(P)$ is totally pathwise differentiable at any $(P,u)\in\mathcal{M}\times\mathcal{U}$ with $\dot{\theta}_{P,u}(s,t)=\dot{\nu}_{P}(s)$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P}^{*}(w),0)$, which will verify the claimed form of the adjoint of the differential operator for the pathwise differentiable primitives in Table 1.
C.3.2   Root-density
@@ C.3.3   Conditional density
The map $\nu_{1}$ is defined so that $\nu_{1}(\tilde{P})=(\tilde{p}_{Y|X}/p_{Y|X})^{1/2}$. Lemma S8 in Luedtke and Chung (2024) shows that $\nu_{1}$ is pathwise differentiable at $P$ with $\dot{\nu}_{1,P}(s)(z)=\frac{1}{2}\Pi_{P}(s)(z)$. The adjoint of this map is $\dot{\nu}_{1,P}^{*}(w)(z)=\frac{1}{2}\{w(z)-E_{P}[w(Z)\,|\,X=x]\}$. The map $\zeta_{1}$ is the pointwise operation $\zeta_{1}(u)(z)=u^{2}(z)p_{Y|X}(z)$. By Appendix C.4.5 and the fact that conditional densities of distributions in $\mathcal{M}$ are uniformly bounded, this map is Hadamard differentiable with $\dot{\zeta}_{1,u}(t)=2tup_{Y|X}$ and $\dot{\zeta}_{1,u}^{*}(w)=2wup_{Y|X}$. The map $\zeta_{2}$ is the change of measure studied in Appendix C.4.9 where, in the notation of that appendix, $\lambda_{1}=P$ and $\lambda_{2}=Q$. By results in that appendix, the fact that $\|dQ/dP\|_{L^{\infty}(P)}<\infty$ implies this map is Hadamard differentiable with $\dot{\zeta}_{2,u}(t)=t$ and $\dot{\zeta}_{2,u}^{*}(w)=\frac{dQ}{dP}w$. Theorem 1 from this work (see also Theorem 3.1 in van der Vaart, 1991a) then shows that $\nu=\zeta_{2}\circ\zeta_{1}\circ\nu_{1}$ is Hadamard differentiable with $\dot{\nu}_{P}^{*}(w)=\dot{\nu}_{1,P}^{*}\circ\dot{\zeta}_{1,\nu_{1}(P)}^{*}\circ\dot{\zeta}_{2,\zeta_{1}\circ\nu_{1}(P)}^{*}(w)$, and properties of adjoints of compositions show that $\dot{\nu}_{P}(s)=\dot{\zeta}_{2,\zeta_{1}\circ\nu_{1}(P)}\circ\dot{\zeta}_{1,\nu_{1}(P)}\circ\dot{\nu}_{1,P}(s)$. Plugging in the provided forms of the differential operators and their adjoints gives (S37).
C.3.4   Dose-response function
@@ C.3.5   Counterfactual density
Under causal conditions, this parameter corresponds to the counterfactual density that would be observed if everyone received treatment $A=1$ (Kennedy et al., 2021). Letting $s_{Y\,|\,A,X}$ and $s_{X}$ be as defined in Appendix C.3.4, the results in Luedtke and Chung (2024) show $\nu$ is pathwise differentiable with local parameter
@@
C.4   Maps that only depend on their Hilbert-valued argument
C.4.1   General case: Hadamard differentiable map
@@
In the remainder of Appendix C.4, we establish the Hadamard differentiability of maps $\zeta:\mathcal{U}\rightarrow\mathcal{W}$. Applying the general results of this Appendix C.4.1 then shows that the parameter $\theta$ with $\theta(P,u)=\zeta(u)$ is totally pathwise differentiable at any $(P,u)\in\mathcal{M}\times\mathcal{U}$ with $\dot{\theta}_{P,u}(s,t)=\dot{\zeta}_{u}(t)$ and $\dot{\theta}_{P,u}^{*}(w)=(0,\dot{\zeta}_{u}^{*}(w))$, which will verify the claimed form of the adjoint of the differential operator for the Hadamard differentiable primitives in Table 1.
C.4.2   Inner product
@@
Importantly, the Hilbert space $\mathcal{T}=\mathcal{R}\oplus\mathcal{R}$ is not ambient for this primitive: the evaluation of this primitive inherently depends on the inner product in $\mathcal{R}$. Hence, when used in Algorithm 1, this primitive does not allow $\mathcal{R}$ to depend on the input to the algorithm, $P$. For example, consider the case where $\mathcal{R}=L^{2}(P)$. In this case, a $P$-dependent inner product primitive would take the form $\theta(P,(u_{1},u_{2}))=\langle u_{1},u_{2}\rangle_{L^{2}(P)}$, which depends nontrivially on its distribution-valued argument and therefore should be studied as in Appendix C.2. For the particular case where $\mathcal{R}=L^{2}(P)$, the $P$-dependent inner product primitive is a special case of the multifold mean studied in Appendix C.2.2, and so is totally pathwise differentiable under conditions given there.
C.4.3   Squared norm
@@
Like the inner product primitive in Appendix C.4.2, the Hilbert space $\mathcal{T}$ is not ambient for this primitive: evaluating $\|\cdot\|_{\mathcal{T}}^{2}$ inherently depends on $\mathcal{T}$. Hence, when used in Algorithm 1, this inner product primitive does not allow $\mathcal{T}$ to depend on the input to the algorithm $P$.
C.4.4   Differentiable functions